# The Evolution of Dataset Distillation: Toward Scalable and Generalizable Solutions

### v2

@@ 2.1.2 Trajectory Matching
To tackle scalability challenges with large-scale datasets, Cui et al. [24] developed Trajectory Matching with Soft Label Assignment (TESLA). TESLA refines the MTT loss function by decomposing gradient computations into sequential batches, eliminating the need to store full computational graphs in memory. Additionally, TESLA incorporates a soft label assignment strategy, leveraging pre-trained teacher models to generate soft labels for synthetic data, which significantly boosts performance when scaling to datasets with a large number of classes, such as ImageNet-1K. Another innovation in this direction is the Automatic Training Trajectories (ATT) method introduced by Liu et al. [25]. Unlike most trajectory matching approaches that use fixed trajectory lengths, ATT dynamically adjusts trajectory lengths during the distillation process. By employing a minimum distance strategy to select optimal matching trajectories, ATT effectively reduces mismatching errors and ensures precise alignment of synthetic datasets.
@@ 2.1.3 Distribution Matching
In the third area, recent approaches utilized higher-order distribution matching, a crucial advancement beyond traditional mean distribution alignment. Zhang et al. [32] pioneered embedding representations into a Reproducing Kernel Hilbert Space [33], enabling alignment across multiple moments. Wei et al. [34] complemented this with Latent Quantile Matching (LQM), aligning synthetic data with optimal quantiles determined via statistical tests. From a geometric perspective, Liu et al. [35] leveraged the Wasserstein distance [36], capturing first and second-order statistics for distribution alignment. Similarly, Deng et al. [37] proposed class centralization and covariance matching, incorporating second-order information to tighten intra-class clustering while maintaining inter-class separation.
Stage Summary Collectively, these methods highlight the transition from simple mean-matching techniques to statistically and geometrically informed alignment strategies. However, as shown in Table II, distribution matching methods often lag behind gradient matching approaches in accuracy on large-scale datasets, revealing fundamental limitations in capturing dynamic training behaviors. To bridge this gap, future research could explore hybrid approaches that integrate distribution and gradient matching, develop adaptive higher-order constraints that evolve during optimization, and investigate causal relationships within feature distributions to gain deeper insights into the effectiveness of different matching strategies.
@@ 2.1.4 Latent and Frequency Space Methods
Operating in latent spaces has emerged as a promising direction for efficient distillation due to its ability to reduce dimensionality while preserving essential information. Duan et al. [38] proposed LatentDD, which transitions from the pixel space to the latent space using a pretrained autoencoder. By condensing datasets into compact latent representations, LatentDD substantially reduces computation and memory requirements, enabling efficient dataset distillation without compromising performance. Building on this shift to latent spaces, George et al. [39] proposed GlaD, which leverages the latent spaces of generative models like StyleGAN-XL [40]. By encoding high-dimensional images into low-dimensional feature vectors, GlaD not only reduces computational cost but also enhances generalization ability.
Complementary to latent space approaches, frequency-domain methods offer another efficient avenue for dataset distillation. Shin et al. [41] introduced Frequency Domain-based Dataset Distillation (FreD), a method that operates in the frequency domain rather than the spatial domain. FreD optimizes a selective subset of important frequency dimensions based on explained variance ratios, avoiding the need to process the full dimensionality of the spatial domain. This approach drastically lowers the memory budget per synthetic instance while preserving critical information. Building upon frequency-domain techniques, Yang et al. [42] introduced Neural Spectral Decomposition (NSD), leveraging the low-rank structure of datasets to decompose them into spectral tensors and kernel matrices, enhancing learning efficiency and integrating seamlessly with frameworks like MTT. Moving beyond traditional representations, Shin et al. [43] introduces Distilling Dataset into Neural Field (DDiF), a novel parameterization framework that compresses large datasets into synthetic neural fields under limited storage budgets, offering superior performance on various datasets.
@@ 2.1.5 Plug-and-Play Approaches
Recent advancements in dataset distillation explored plug-and-play techniques that seamlessly integrate into existing methods, enhancing both optimization efficiency and generalization. Unlike approaches that introduce entirely new optimization paradigms, these methods serve as modular enhancements, improving adaptability and scalability. However, it is important to note that current plug-and-play techniques are primarily designed for matching-based methods, such as gradient matching and trajectory matching.
@@ 2.2.1.2   Diffusion-based Methods
Diffusion models, which are renowned for their superior image quality compared to GANs, have emerged as a powerful alternative. Gu et al. [60] pioneered a diffusion-based framework leveraging a ‘minimax’ criterion to balance representativeness and diversity in surrogate datasets, ensuring both coverage and variability of training data. Similarly, Su et al. [61] introduced Dataset Distillation via Disentangled Diffusion Model (D4M), which combines latent diffusion models with prototype learning. This approach enhances distillation efficiency and cross-architecture generalization while employing a novel training-time matching strategy. By eliminating architecture-specific optimization during image synthesis, D4M achieves high-resolution image generation, representing a substantial leap forward in dataset distillation methodologies.
@@
where $T$ denotes the total number of diffusion steps, $t$ represents the current timestep in the backward process, $\mu_{\theta}$ is the parameterized mean predictor network with learnable parameters $\theta$, and ${z}_{T}$ is the initial noisy state. The linear interpolation coefficients $(1-\frac{t}{T})$ and $\frac{t}{T}$ control the balance between the mean prediction and initial state, providing enhanced gradient flow during training without compromising the generation quality. This framework demonstrates superior performance for high-resolution tasks while integrating seamlessly with existing algorithms. Taking a different approach to latent space optimization, Qin et al. [66] proposed Distributional Dataset Distillation (D3) and applied it into the federated learning setting [67]. In their method, D3 represents classes as Gaussian distributions in latent space, optimizing the parameters to generate images using a decoder network. To guide the distilled data generation process, in [68], Chen et al. introduced Influence-Guided Diffusion (IGD) that uses a latent diffusion model based on a Diffusion Transformer [69] to generate synthetic training datasets. They designed an influence guidance function to steer diffusions towards generating training-effective data and a deviation guidance function to enhance diversity, enabling the method to significantly improve performance across various datasets, particularly on high-resolution ImageNet-1K benchmarks.
@@ 2.3.1 Selective Dataset Distillation
Selective Pixel Enhancement Effectively identifying and enhancing discriminative pixels is crucial for dataset quality, especially when addressing inter-class variability and non-discriminative regions. Recent efforts have refined pixel-wise feature representation to emphasize informative pixels while suppressing redundant or misleading patterns. Wang et al. [86] proposed the Emphasizing Discriminative Features (EDF) framework, which leverages Common Pattern Dropout (CPD) to filter out non-salient pixels and Discriminative Area Enhancement (DAE) to intensify attention on essential regions using Grad-CAM [87] activation maps. By adaptively selecting and enhancing discriminative regions, EDF ensures synthesized datasets prioritize high-salience regions, leading to improved generalization ability.
@@
Prioritization methods focus on dynamically ranking and weighting samples during training to emphasize their relative importance. Wang et al. [94] introduced Sample Difficulty Correction (SDC), which prioritizes simpler samples based on gradient norms. To adapt dataset difficulty dynamically, Lee et al. [95] proposed SelMatch, which employs selection-based initialization and partial updates to fine-tune synthetic datasets. By tailoring the training process to specific architectures and datasets, SelMatch improves generalization and convergence. Chen et al. [96] extended prioritization by introducing an adversarial prediction matching framework. This approach leverages teacher-student disagreements to identify informative samples, enhancing robustness across diverse architectures. Tukan et al. [97] proposed adaptive sampling and initialization strategies. Focusing on data prioritization, Li et al. [98] proposed Prioritize Alignment in Dataset Distillation (PAD), improving distillation effectiveness through two key strategies: Data Filtering, which dynamically selects high-complexity samples based on EL2N scores [99] to retain the most informative data, and Layer Confinement, which restricts the distillation process to deeper model layers, refining the prioritization of critical features while reducing noise from low-level representations.
@@ 2.3.5 Extreme Compression Techniques
Extreme compression techniques address storage constraints by drastically reducing dataset sizes while preserving training effectiveness. Shul et al. [110] introduced Poster Dataset Distillation (PoDD), which compresses entire datasets into a single poster image, achieving extreme compactness with less than one image per class. By leveraging semantic class ordering and efficient label management, PoDD balances high compression with robust training performance, making it well-suited for resource-constrained applications. Future research could explore expanding extreme compression strategies to large-scale datasets while improving adaptability across different model architectures.
@@ 2.4.1 Addressing Non-IID Challenges
Out-Of-Distributions In open-world scenarios, models must reliably identify and handle data samples that deviate from their training distribution. Traditional dataset distillation methods overlook this, focusing on IID data. To address this gap, Ma et al. proposed Trustworthy Dataset Distillation (TrustDD) [116], which integrates OOD detection into the distillation process. By generating synthetic outliers through Pseudo-Outlier Exposure (POE), TrustDD eliminates the need for curated OOD datasets, enabling robust OOD detection without compromising in-distribution (InD) accuracy or computational efficiency.
@@ 2.4.2 Addressing Non-centralized Challenges
Privacy Preservation Privacy concerns in FL have driven the integration of dataset distillation techniques that limit information sharing. FedDGM (Federated Dataset Distillation with Deep Generative Models) [134] utilized server-side pre-trained generators to distill datasets in latent space, avoiding the need to share raw data or updates while maintaining high performance. QUICKDROP [135] extended this concept by embedding dataset distillation into federated unlearning, reducing computational costs without sacrificing privacy or performance. Xu et al. [136] applied the principle of least privilege (PoLP), where clients share only essential knowledge through local-global dataset distillation, further safeguarding sensitive information.
@@ 3.1 Temporal Domain
Video Domain Wang et al. [157] conducted the first systematic study on video dataset distillation, classifying temporal compression methods into four dimensions: synthetic frames, real frames, segments, and interpolation algorithms. They proposed a two-stage framework that disentangles static and dynamic information, integrating single-frame static memory with dynamic memory blocks to synthesize realistic video data. Building on this, Chen et al. [158] introduced an efficient temporal condensation strategy incorporating slide-window sampling and adaptive trajectory matching, significantly improving performance on action recognition benchmarks while reducing training costs.

### v3

@@ 2.1.2 Trajectory Matching
To tackle scalability challenges with large-scale datasets, Cui et al. [24] developed Trajectory Matching with Soft Label Assignment (TESLA). TESLA refines the MTT loss function by decomposing gradient computations into sequential batches, eliminating the need to store full computational graphs in memory. Additionally, TESLA incorporates a soft label assignment strategy, leveraging pre-trained teacher models to generate soft labels for synthetic data, which significantly boosts performance when scaling to datasets with a large number of classes, such as ImageNet-1K. Another innovation in this direction is the Automatic Training Trajectories (ATT) method introduced by Liu et al. [25]. Unlike most trajectory matching approaches that use fixed trajectory lengths, ATT dynamically adjusts trajectory lengths during the distillation process. By employing a minimum distance strategy to select optimal matching trajectories, ATT effectively reduces mismatching errors and ensures precise alignment of synthetic datasets. Kong et al. [26] propose MTT-LSS, another trajectory matching-based method. Unlike previous methods emphasizing large-scale datasets, MTT-LSS primarily aims at reducing computational overhead and memory usage through low-rank space sampling. It achieves this by representing synthetic data with shared dimension mappers and low-dimensional basis vectors. Further evaluations on larger-scale datasets could help demonstrate its full potential regarding scalability.
@@ 2.1.3 Distribution Matching
In the third area, recent approaches utilized higher-order distribution matching, a crucial advancement beyond traditional mean distribution alignment. Zhang et al. [33] pioneered embedding representations into a Reproducing Kernel Hilbert Space [34], enabling alignment across multiple moments. Wei et al. [35] complemented this with Latent Quantile Matching (LQM), aligning synthetic data with optimal quantiles determined via statistical tests. From a geometric perspective, Liu et al. [36] leveraged the Wasserstein distance [37], capturing first and second-order statistics for distribution alignment. Similarly, Deng et al. [38] proposed class centralization and covariance matching, incorporating second-order information to tighten intra-class clustering while maintaining inter-class separation. Wang et al. [39] propose a frequency-domain-based method that redefines distribution matching using the Neural Characteristic Function (NCF) and employs a minmax optimization strategy to dynamically learn an optimal discrepancy metric, ensuring a more precise alignment between distilled and real data distributions. The method first maximizes the discrepancy between real and distilled data in a learned frequency-space representation to establish the most discriminative distribution metric. Then, it minimizes this discrepancy by optimizing the distilled data to align with the real data under this learned metric. Unlike traditional methods that operate in pixel or feature space, this approach matches both phase and amplitude information in the frequency domain, enhancing the realism and diversity of distilled data while significantly improving computational efficiency.
Stage Summary Collectively, these methods highlight the transition from simple mean-matching techniques to statistically and geometrically informed alignment strategies. Notably, recent work by Li et al. introduced Hyperbolic Dataset Distillation (HDD) [40], leveraging hyperbolic geometry to explicitly integrate hierarchical relationships into distribution matching, significantly improving representational effectiveness. However, as shown in Table II, distribution matching methods often lag behind gradient matching approaches in accuracy on large-scale datasets, revealing fundamental limitations in capturing dynamic training behaviors. To bridge this gap, future research could explore hybrid approaches that integrate distribution and gradient matching, develop adaptive higher-order constraints that evolve during optimization, and investigate causal relationships within feature distributions to gain deeper insights into the effectiveness of different matching strategies.
@@ 2.1.4 Latent and Frequency Space Methods
Operating in latent spaces has emerged as a promising direction for efficient distillation due to its ability to reduce dimensionality while preserving essential information. Duan et al. [41] proposed LatentDD, which transitions from the pixel space to the latent space using a pretrained autoencoder. By condensing datasets into compact latent representations, LatentDD substantially reduces computation and memory requirements, enabling efficient dataset distillation without compromising performance. Building on this shift to latent spaces, George et al. [42] proposed GlaD, which leverages the latent spaces of generative models like StyleGAN-XL [43]. By encoding high-dimensional images into low-dimensional feature vectors, GlaD not only reduces computational cost but also enhances generalization ability. Further extending latent space methods, Li et al. proposed SLFD [44], explicitly modeling spatial correlations and uncertainty within latent features using a low-rank multivariate normal distribution, effectively capturing structured uncertainty for improved dataset distillation.
Complementary to latent space approaches, frequency-domain methods offer another efficient avenue for dataset distillation. Shin et al. [45] introduced Frequency Domain-based Dataset Distillation (FreD), a method that operates in the frequency domain rather than the spatial domain. FreD optimizes a selective subset of important frequency dimensions based on explained variance ratios, avoiding the need to process the full dimensionality of the spatial domain. This approach drastically lowers the memory budget per synthetic instance while preserving critical information. Building upon frequency-domain techniques, Yang et al. [46] introduced Neural Spectral Decomposition (NSD), leveraging the low-rank structure of datasets to decompose them into spectral tensors and kernel matrices, enhancing learning efficiency and integrating seamlessly with frameworks like MTT. Moving beyond traditional representations, Shin et al. [47] introduces Distilling Dataset into Neural Field (DDiF), a novel parameterization framework that compresses large datasets into synthetic neural fields under limited storage budgets, offering superior performance on various datasets. To enhance the flexibility of frequency-based dataset distillation, Bo et al. [48] proposed Curriculum Frequency Matching (CFM), which progressively refines filter parameters throughout the distillation process. Unlike static filtering methods, CFM dynamically transitions between low- and high-frequency components, improving adaptability and generalization.
@@ 2.1.5 Plug-and-Play Approaches
Recent advancements in dataset distillation, such as [49, 50, 51, 52, 53, 54, 55, 56], explored plug-and-play techniques that seamlessly integrate into existing methods, enhancing both optimization efficiency and generalization. Unlike approaches that introduce entirely new optimization paradigms, these methods serve as modular enhancements, improving adaptability and scalability. However, it is important to note that current plug-and-play techniques are primarily designed for matching-based methods, such as gradient matching and trajectory matching.
@@
Similarly, OPTICAL [50] addresses uniform optimization issues by introducing entropy-regularized Optimal Transport matching as a plug-and-play module. Specifically, OPTICAL assigns different contribution weights from real to synthetic samples based on optimal transport matching of feature distances, allowing synthetic datasets to better preserve fine-grained geometric details. This leads to consistent improvements in generalization performance without requiring significant changes to existing gradient or trajectory matching frameworks.
@@ 2.2.1.2   Diffusion-based Methods
Diffusion models, which are renowned for their superior image quality compared to GANs, have emerged as a powerful alternative. Gu et al. [67] pioneered a diffusion-based framework leveraging a ‘minimax’ criterion to balance representativeness and diversity in surrogate datasets, ensuring both coverage and variability of training data. Similarly, Su et al. [68] introduced Dataset Distillation via Disentangled Diffusion Model (D4M), which combines latent diffusion models with prototype learning. This approach enhances distillation efficiency and cross-architecture generalization while employing a novel training-time matching strategy. By eliminating architecture-specific optimization during image synthesis, D4M achieves high-resolution image generation, representing a substantial leap forward in dataset distillation methodologies. To address the misalignment between generation and evaluation objectives, CaO2 [69] introduces a two-stage framework that first selects class-consistent samples using a lightweight classifier, then refines their latent codes to improve conditional alignment. This design effectively mitigates misalignment between generation and evaluation objectives, leading to improved distillation quality across diverse settings. To enhance semantic control and mitigate object inconsistency in generated images, Zou et al. [70] proposed a vision-language distillation framework that introduces category-wise text prototypes alongside image prototypes to guide diffusion-based image generation. By leveraging large language models to produce cluster-specific descriptions and integrating them into the sampling process, their method improves semantic coherence, reduces object omissions, and achieves superior accuracy across various benchmarks.
@@
where $T$ denotes the total number of diffusion steps, $t$ represents the current timestep in the backward process, $\mu_{\theta}$ is the parameterized mean predictor network with learnable parameters $\theta$, and ${z}_{T}$ is the initial noisy state. The linear interpolation coefficients $(1-\frac{t}{T})$ and $\frac{t}{T}$ control the balance between the mean prediction and initial state, providing enhanced gradient flow during training without compromising the generation quality. This framework demonstrates superior performance for high-resolution tasks while integrating seamlessly with existing algorithms. Recently, Zhao et al. proposed D³HR [75], a diffusion-based dataset distillation framework leveraging DDIM inversion to map complex latent distributions into a structured Gaussian space for better distribution matching. Additionally, their novel group sampling strategy selects representative subsets based on statistical similarity, significantly improving accuracy, stability, and cross-architecture generalization compared to prior methods. Taking a different approach to latent space optimization, Qin et al. [76] proposed Distributional Dataset Distillation (D3) and applied it into the federated learning setting [77]. In their method, D3 represents classes as Gaussian distributions in latent space, optimizing the parameters to generate images using a decoder network. To guide the distilled data generation process, in [78], Chen et al. introduced Influence-Guided Diffusion (IGD) that uses a latent diffusion model based on a Diffusion Transformer [79] to generate synthetic training datasets. They designed an influence guidance function to steer diffusions towards generating training-effective data and a deviation guidance function to enhance diversity, enabling the method to significantly improve performance across various datasets, particularly on high-resolution ImageNet-1K benchmarks. Chan et al. [80] proposed MGD3, a training-free dataset distillation framework that leverages pre-trained diffusion models and injects semantic mode guidance directly into the sampling process to generate diverse and representative samples—without any model parameter updates.
@@ 2.3.1 Selective Dataset Distillation
Selective Pixel Enhancement Effectively identifying and enhancing discriminative pixels is crucial for dataset quality, especially when addressing inter-class variability and non-discriminative regions. Recent efforts have refined pixel-wise feature representation to emphasize informative pixels while suppressing redundant or misleading patterns. Wang et al. [97] proposed the Emphasizing Discriminative Features (EDF) framework, which leverages Common Pattern Dropout (CPD) to filter out non-salient pixels and Discriminative Area Enhancement (DAE) to intensify attention on essential regions using Grad-CAM [98] activation maps. By adaptively selecting and enhancing discriminative regions, EDF ensures synthesized datasets prioritize high-salience regions, leading to improved generalization ability. Extending this idea of emphasizing discriminative pixels, Tran et al. [99] proposed a region-based refinement strategy, which also uses CAM to precisely identify critical regions and preserve essential instance-specific details. In their method, non-critical regions are selectively refined to effectively incorporate broader class-general information into the distilled dataset.
@@
Prioritization methods focus on dynamically ranking and weighting samples during training to emphasize their relative importance. Wang et al. [106] introduced Sample Difficulty Correction (SDC), which prioritizes simpler samples based on gradient norms. To adapt dataset difficulty dynamically, Lee et al. [107] proposed SelMatch, which employs selection-based initialization and partial updates to fine-tune synthetic datasets. By tailoring the training process to specific architectures and datasets, SelMatch improves generalization and convergence. Extending the idea of curriculum-based sample prioritization, Chen et al. [108] proposed a curriculum-based coarse-to-fine selection method to enhance high-IPC dataset distillation by adaptively selecting real training samples based on their difficulty. This method progressively excludes easy examples and carefully integrates challenging yet manageable instances, significantly improving the representativeness and generalization capability of distilled datasets. Chen et al. [109] further extended prioritization by introducing an adversarial prediction matching framework. This approach leverages teacher-student disagreements to identify informative samples, enhancing robustness across diverse architectures. Tukan et al. [110] proposed adaptive sampling and initialization strategies. Focusing on data prioritization, Li et al. [111] proposed Prioritize Alignment in Dataset Distillation (PAD), improving distillation effectiveness through two key strategies: Data Filtering, which dynamically selects high-complexity samples based on EL2N scores [112] to retain the most informative data, and Layer Confinement, which restricts the distillation process to deeper model layers, refining the prioritization of critical features while reducing noise from low-level representations.
@@ 2.3.5 Extreme Compression Techniques
Extreme compression techniques address storage constraints by drastically reducing dataset sizes while preserving training effectiveness. Shul et al. [123] introduced Poster Dataset Distillation (PoDD), which compresses entire datasets into a single poster image, achieving extreme compactness with less than one image per class. By leveraging semantic class ordering and efficient label management, PoDD balances high compression with robust training performance, making it well-suited for resource-constrained applications. Li et al. [124] propose DATM-CLR to address the semantic degradation of synthetic data under extreme compression, e.g., IPC=1. By integrating supervised contrastive learning into the image synthesis stage, DATM-CLR enhances intra-class coherence and inter-class separability, achieving a significant accuracy gain over DATM on CIFAR-10 and consistent improvements across CIFAR-100 and Tiny-ImageNet.
@@ 2.4.1 Addressing Non-IID Challenges
Out-Of-Distribution and Cross-Domain Generalization In open-world scenarios, models must reliably identify and handle data samples that deviate from their training distribution. Traditional dataset distillation methods overlook this, typically focusing only on IID data. To explicitly address the OOD detection problem within dataset distillation, Ma et al. introduced Trustworthy Dataset Distillation (TrustDD) [130], integrating synthetic outlier generation via Pseudo-Outlier Exposure into the distillation process. TrustDD thereby eliminates dependence on curated OOD datasets, enabling robust OOD detection without sacrificing in-distribution accuracy or computational efficiency. Moving beyond single-domain scenarios, Choi et al. [131] addressed a related but distinct challenge: multi-domain dataset condensation. They proposed a Domain-Aware Module, employing frequency-based pseudo-domain labeling to automatically partition datasets into pseudo-domains, and learnable spatial domain masks to explicitly encode domain-specific features, significantly improving the synthesized dataset’s generalization across diverse visual domains.
@@ 2.4.2 Addressing Non-centralized Challenges
Privacy Preservation Privacy concerns in FL have driven the integration of dataset distillation techniques that limit information sharing. FedDGM [149] utilizes pre-trained generators on the server to distill datasets into latent space, effectively preserving privacy by avoiding direct sharing of sensitive data. QUICKDROP [150] extended this concept by embedding dataset distillation into federated unlearning, reducing computational costs without sacrificing privacy or performance. Xu et al. [151] applied the principle of least privilege (PoLP), where clients share only essential knowledge through local-global dataset distillation, further safeguarding sensitive information. Moreover, FedWSIDD [152] specifically addresses medical imaging scenarios, proposing a federated learning framework that exchanges synthetic slides instead of real patient data or model parameters, facilitating secure and efficient collaboration among diverse medical institutions.
@@ 3.1 Temporal Domain
Video Domain Wang et al. [173] conducted the first systematic study on video dataset distillation, classifying temporal compression methods into four dimensions: synthetic frames, real frames, segments, and interpolation algorithms. They proposed a two-stage framework that disentangles static and dynamic information, integrating single-frame static memory with dynamic memory blocks to synthesize realistic video data. Building on this, Chen et al. [174] introduced an efficient temporal condensation strategy incorporating slide-window sampling and adaptive trajectory matching, significantly improving performance on action recognition benchmarks while reducing training costs. Most recently, Li et al. [175] proposed the first latent-space video dataset distillation framework, combining VAE-based video encoding with high-order singular value decomposition (HOSVD) for compression. Leveraging latent representations introduced significant theoretical advancements, though practical implications for video quality merit further exploration. Furthermore, Li et al. [176] proposed a novel video dataset distillation framework leveraging a pretrained video diffusion model, Latte [177], to efficiently generate high-quality synthetic videos. Subsequently, two innovative video selection strategies were introduced: (1) VST-UNet, a 4D spatio-temporal U-Net architecture optimized via cross-entropy, diversity, and representativeness losses, and (2) TAC-DT, a training-free, temporal-aware clustering method utilizing VideoMAE embeddings and hierarchical BIRCH clustering. Employing diffusion models and advanced selection strategies demonstrated substantial benefits, marking a promising future direction in the field.